OpenStudy (anonymous):

How do you factor 4x2-81? (That's 4 x squared)

OpenStudy (aroub):

(2x-9)(2x+9)

OpenStudy (anonymous):

(2x+9)(2x-9)

OpenStudy (anonymous):

4(x^2 -81/4) 4(x- 9/2)(x+9/2) above answers are right too

OpenStudy (anonymous):

its the difference of squares formula a^2-b^2=(a-b)(a+b) here a=2x b=9